OpenStudy (anonymous):

Solve the following system of equations x+3y=7 (1) x=9-3y (2) Solution should be an ordered pair.

OpenStudy (bahrom7893):

no solutions..

OpenStudy (bahrom7893):

9-3y+3y = 7 9 = 7 thats false

OpenStudy (anonymous):

Grrrrrr...

OpenStudy (anonymous):

Well that's better, but still.

OpenStudy (anonymous):

So you put the two equations together?

OpenStudy (bahrom7893):

yea.. substitute for x in the 1st eqn..

OpenStudy (anonymous):

You can alternatively add the equations.. x + 3y = 7 + -3y + 9 = x ============ x + 9 = 7 + x So we have: x + 9 = x + 7 Which is impossible.

OpenStudy (anonymous):

ahhhh that makes sense